Complete the balance sheet and sales information below for Iceberg Industries using the following financial data: Debt ratio: 50% Quick Ratio: .80x Total Assets Turnover: 2.0x Days Sales Outstanding:36.0 days Gross Profit Margin on Sales (Sales-Cost of Goods Sold)/Sales-25% Inventory Turnover Ratio: 5.0x Balance Sheet: Cash Select ] . Accounts Receivable I Select ] Inventories Select . Fixed Assets [ Select ] . Total Assets $300,000 . Accounts Payable . Long Term Debt $70,000 . Common Stock ISelect . Retained Earnings $97,500

. Retained Earnings $97,500 Total Liabilities and Equity Select) . Sales Select ] Cost of Goods Sold [Select ]
Answer #1
Total Assets turnover: 2.0 x
Total assets: 300000
Total assets trunover = Sales/ Total assets
2 = Sales / 300000
Sales =$ 600000
Gross profit ratio as % of sales =25%
Gross profit = 600000*25% = 150000
Cost of goods sold = Sales - Gross Profit
600000- 150000 = 450000
Inventory turnover ratio = Cost of goods sold / Inventory
5 = 450000 / Inventory
Inventory = 90000
Days sales outstanding = 36
Days sales outstanding = Number of days in year / sales * accounts receivable
36 = 360 / 600000 * Accounts receivable
Accounts receivable =60000
Debt ratio = 50%
Debt rtio = Total debts / Total assets
0.50 = Total debts / 300000
Total debts = 150000
Total debts = Long term debts + Current liabilities
150000 =70000 + Current liabilities
Current liabilities (Accounts payable) = 80000
Quick Ratio = 0.80
Quick ratio = Liquid assets / Current liabilities
0.80 = Liquid assets / 80000
Liquid assets = 64000
Liquid assets = Cash + Accounts receivable
64000 = Cash + 60000
Cash = 4000
Total assets = Cash +Accounts receivable + Inventory + Fixed assets
300000 = 4000+60000+90000+ Fixed assets
Fixed assets = 146000
Total assets = Total liabilities + Stockholder's equity
300000 = 150000 + 150000
Stockholder's equity = 150000
Stockholder's equity = Common Stock + Retained earnings
150000 = Common Stock + 97500
Common stock = 52500
Final Answer is:
Balance Sheet:
Cash 4000
Accounts receivbale 60000
Inventory 90000
Fixed assets 146000
Total assets 300000
Accounts payable 80000
Long term debts 70000
Common Stock 52500
Retained earnings 97500
Total liabilities & Equity 300000
Sales 600000
Cost of goods sold 450000